OpenStudy (anonymous):

the solution to 64^(4-x)=4^(2x)

jimthompson5910 (jim_thompson5910):

Rewrite 64 as 4^3 to get (4^3)^(4-x) = 4^(2x) Then multiply exponents to get 4^(3(4-x)) = 4^(2x) Since the bases are equal, the exponents must be equal, so 3(4-x) = 2x Solve that for x to get your answer.

OpenStudy (anonymous):

is it x=12/5

jimthompson5910 (jim_thompson5910):

you got it, nice job

OpenStudy (anonymous):

yay! :)